What is t| | z 10.00 0.01 0.02 0.03 0.04 0.05 0.06 0.07 0.08 0.09 1.8 0.036 0.035 0.034 0.034 0.033 0.032 0.031 0.031 0.030 0.029 1.7|he p-value of a two-tailed one-mean hypothesis test, with a test statistic of z0=−1.59? (Do not round your answer; compute your answer using a value from the table below.)
Answer #1

for given two tailed one-mean hypothesis test

p value of the test =2*P(Z<-1.59) =2*(0.056)=0.112